# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v2

@@ A.1 Related Works
However, as we noted in the Section 1, one can do a greedy approximation of the optimal action from the historical data using a weak demonstrator and a neural network policy (Finn et al., 2017; Rothfuss et al., 2018). Moreover, the in-context framework generally is more agnostic where it learns the policy of the demonstrator (Duan et al., 2016; Wang et al., 2016; Mishra et al., 2017). Note that both DPT-greedy and PreDeToR are different than algorithmic distillation (Laskin et al., 2022; Lu et al., 2023) as they do not distill an existing RL algorithm. moreover, in contrast to DPT-greedy which is trained to predict the optimal action, the PreDeToR is trained to predict the reward for each of the actions. This enables the PreDeToR (similar to DPT-greedy) to show to potentially emergent online and offline strategies at test time that automatically align with the task structure, resembling posterior sampling.
As we discussed in the Section 1, in decision-making, RL, and imitation learning the transformer models are trained using autoregressive action prediction (Yang et al., 2023). Similar methods have also been used in Large language models (Vaswani et al., 2017; Roberts et al., 2019). One of the more notable examples is the Decision Transformers (abbreviated as DT) which utilizes a transformer to autoregressively model sequences of actions from offline experience data, conditioned on the achieved return (Chen et al., 2021; Janner et al., 2021). This approach has also been shown to be effective for multi-task settings (Lee et al., 2022), and multi-task imitation learning with transformers (Reed et al., 2022; Brohan et al., 2022; Shafiullah et al., 2022). However, the DT methods are not known to improve upon their in-context data, which is the main thrust of this paper (Brandfonbrener et al., 2022; Yang et al., 2022b).
@@ A.2.2 Details of Baselines
(1) Thomp: This baseline is the stochastic $A$-action bandit Thompson Sampling algorithm from Thompson (1933); Agrawal & Goyal (2012); Russo et al. (2018); Zhu & Tan (2020). We briefly describe the algorithm below: At every round $t$ and each action $a$, Thomp  samples $\gamma_{m,t}(a)\sim\mathcal{N}(\widehat{\mu}_{m,t-1}(a),\sigma^{2}/N_{m,t-1}(a))$, where $N_{m,t-1}(a)$ is the number of times the action $a$ has been selected till $t-1$, and $\widehat{\mu}_{m,t-1}(a)=\frac{\sum_{s=1}^{t-1}\widehat{r}_{m,s}\mathbf{1}(I_{s}=a)}{N_{m,t-1}(a)}$ is the empirical mean. Then the action selected at round $t$ is $I_{t}=\mathop{\rm arg\,max}_{a}\gamma_{m,t}(a)$. Observe that Thomp is not a deterministic algorithm like UCB (Auer et al., 2002). So we choose Thomp as the weak demonstrator $\pi^{w}$ because it is more exploratory than UCB and also chooses the optimal action, $a_{m,*}$, a sufficiently large number of times. Thomp is a weak demonstrator as it does not have access to the feature set $\mathcal{X}$ for any task $m$.
@@ A.3 Empirical Study: Comparison against K-armed bandits and DPT
Baselines: In the K-armed bandits We implement the same baselines discussed in Section 4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LinUCB. In the linear and non-linear setting, we compare against DPT instead of DPT-greedy.
Settings: In the K-armed bandit setting we consider $d=6$, and the arms as canonical vectors $\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{6}$. For each task $m$, we choose the hidden parameter $\bm{\theta}_{m,*}$ similar to the linear bandit setting discussed in Section 5. Note that this results in a K-armed bandit setting. For the linear and non-linear setting comparison, we use the same setting as Section 5, and 4.
@@
Experimental Result: We observe these outcomes in Figure 7. In Figure 7(a) the demonstrator $\pi^{w}$ is the Thomp algorithm. Note that there is no structure across arms now, and sampling one arm gives no information about other arms in a task. We observe that PreDeToR-$\tau$ performs similarly to the demonstrator Thomp, and also shows that incorporating exploration is a sound technique. Also, AD performs similarly to the demonstrator Thomp. Both DPT-greedy and PreDeToR fail to learn the latent structure across the tasks and therefore do not learn any exploration strategy.
In Figure 7(b) we show the linear bandit setting discussed in Section A.2. We observe that PreDeToR (-$\tau$) matches the performance of DPT, and LinUCB. Note that DPT has access to the optimal action per task, and LinUCB is the optimal oracle algorithm that leverages the structure information.
In Figure 7(c) we show the non-linear bandit setting discussed in Section A.2. Again we observe that PreDeToR (-$\tau$) matches the performance of DPT and has lower cumulative regret than AD and LinUCB which fails to perform well in this non-linear setting due to its algorithmic design.
@@ A.4 Empirical Study: Bilinear Bandits
Baselines: We again implement the same baselines discussed in Section 4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, and Thomp. Note that we do not implement the LinUCB and MLinGreedy for the bilinear bandit setting. However, we now implement the LowOFUL (Jun et al., 2019) which is optimal in the bilinear bandit setting.
@@
$\widetilde{\mathcal{A}}_{m}$ =$=\left\{\left[\operatorname{vec}\left(\mathbf{x}_{m,1:\kappa}\mathbf{z}_{m,1:\kappa}^{\top}\right);\operatorname{vec}\left(\mathbf{x}_{m,\kappa+1:d_{1}}\mathbf{z}_{m,1:\kappa}^{\top}\right);\operatorname{vec}\left(\mathbf{x}_{m,1:\kappa}\mathbf{z}_{m,\kappa+1:d_{2}}^{\top}\right);\right.\right.$
@@
(a) Rank $1$ $\mathbf{\Theta}_{m,*}$
(b) Rank $2$ $\mathbf{\Theta}_{m,*}$
@@
Experimental Result: We observe these outcomes in Figure 8. In Figure 8(a) we experiment with rank $1$ hidden parameter $\mathbf{\Theta}_{m,*}$ and set horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=30$, and $d=5$. In Figure 8(b) we experiment with rank $2$ hidden parameter $\mathbf{\Theta}_{m,*}$ and set horizon $n=20$, $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR has lower cumulative regret than DPT-greedy, AD and Thomp. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Note that for this small horizon setting the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. In contrast PreDeToR learns the correlation of rewards across tasks and can perform well. Observe from Figure 8(a), and 8(b) that PreDeToR has lower regret than Thomp and matches LowOFUL. Also, in this low-data regime it is not enough for LowOFUL to learn the underlying $\mathbf{\Theta}_{m,*}$ with high precision. Hence, PreDeToR also has slightly lower regret than LowOFUL. Note that the main objective of AD is to match the performance of its demonstrator. Most importantly it shows that PreDeToR can exploit the underlying latent structure and reward correlation better than DPT-greedy, and AD.
@@ A.5 Empirical Study: Latent Bandits
Baselines: We again implement the same baselines discussed in Section 4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LowOFUL. However, we now implement a special LowOFUL (stated in Section A.4) which has knowledge of the shared latent parameters $\mathbf{U}$, and $\mathbf{V}$. We call this the LowOFUL (oracle) algorithm. Therefore LowOFUL (oracle) has knowledge of the problem parameters in the latent bandit setting and hence the name. Again note that we do not implement the LinUCB and MLinGreedy for the latent bandit setting.
@@
(a) Rank $1$ $\mathbf{Z}_{m,*}$
(b) Rank $2$ $\mathbf{Z}_{m,*}$
(c) Rank $3$ $\mathbf{Z}_{m,*}$
@@
Experimental Result: We observe these outcomes in Figure 9. In Figure 9(a) we experiment with rank $1$ hidden parameter $\bm{\theta}_{m,*}\bm{\theta}_{m,*}^{\top}$ and latent parameters $\mathbf{u}\mathbf{v}^{\top}$ shared across the tasks and set horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=30$, and $d=5$. In Figure 9(b) we experiment with rank $2$ hidden parameter $\mathbf{\Theta}_{m,*}$, and latent parameters $\mathbf{U}\mathbf{V}^{\top}$ and set horizon $n=20$, $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. In Figure 9(c) we experiment with rank $3$ hidden parameter $\mathbf{\Theta}_{m,*}$, and latent parameters $\mathbf{U}\mathbf{V}^{\top}$ and set horizon $n=20$, $M_{\text{pre}}=300000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD and Thomp. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Note that for this small horizon setting the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. In contrast PreDeToR (-$\tau$) learns the correlation of rewards across tasks and is able to perform well. Observe from Figure 9(a), 9(b), and 9(c) that PreDeToR has lower regret than Thomp and has regret closer to LowOFUL (oracle)which has access to the problem-dependent parameters. Hence. LowOFUL (oracle) outperforms PreDeToR (-$\tau$) in this setting. This shows that PreDeToR is able to exploit the underlying latent structure and reward correlation better than DPT-greedy, and AD.
@@ A.6 Connection between PreDeToR and Linear Multivariate Gaussian Model
$\mathbf{D}_{A}(i,i)$ =$=\begin{cases}\frac{\sigma^{2}}{N_{n}(a)},\text{ if }N_{n}(a)>0\\ =0,\text{ if }N_{n}(a)=0\end{cases}\vspace*{-1em}$ (6)
where the reward noise being $\sigma^{2}$ sub-Gaussian is known. Finally define the estimated reward covariance matrix $\mathbf{S}_{A}\in\mathbb{R}^{A\times A}$ as $\mathbf{S}_{A}(a,a^{\prime})=\widehat{\mu}_{n}(a)\widehat{\mu}_{n}(a^{\prime})$, where $\widehat{\mu}_{n}(a)$ is the empirical mean of action $a$ estimated from the pretraining data. This matrix captures the reward correlation between the pairs of actions $a,a^{\prime}\in[A]$. Then the posterior average mean estimator $\widehat{\mu}\in\mathbb{R}^{A}$ over all tasks is given by the following lemma. The proof is given in Section B.1.
@@
We hypothesize that the PreDeToR is learning the reward correlation covariance matrix from the training data $\mathcal{H}_{\text{train}}$ and acting greedily on it. To test this hypothesis, we consider the greedy BayesPred algorithm that first estimates $\mathbf{S}_{A}$ from the pretraining data. It then uses the LMMSE estimator in 1 to calculate the posterior mean vector $\widehat{\mu}$, and then selects $I_{t}=\mathop{\rm arg\,max}_{a}\widehat{\mu}(a)$ at each round $t$. Note that BayesPred is a greedy algorithm that always selects the most rewarding action (exploitation) without any exploration of sub-optimal actions. Also the BayesPred is an LMMSE estimator that leverages the linear reward structure and estimates the reward covariance matrix, and therefore can be interpreted as a lower bound to the regret of PreDeToR. The hypothesis that BayesPred is a lower bound to PreDeToR is supported by Figure 10. In Figure 10 the reward covariance matrix for BayesPred is estimated from the $\mathcal{H}_{\text{train}}$ by first running the Thomp ($\pi^{w}$). Observe that the BayesPred has a lower cumulative regret than PreDeToR and almost matches the regret of PreDeToR towards the end of the horizon. Also note that LinUCB has lower cumulative regret towards the end of horizon as it leverages the linear structure and the feature of the actions in selecting the next action.

### v3

@@ 9 Conclusions, Limitations and Future Works
Acknowledgments: Q. Xie is supported in part by National Science Foundation (NSF) grants CNS-1955997 and EPCN-2339794 and EPCN-2432546. J. Hanna is supported in part by NSF (IIS-2410981) and Sandia National Labs through a University Partnership Award.
@@ A.1 Related Works
However, as we noted in the Section˜1, one can do a greedy approximation of the optimal action from the historical data using a weak demonstrator and a neural network policy (Finn et al., 2017; Rothfuss et al., 2018). Moreover, the in-context framework generally is more agnostic where it learns the policy of the demonstrator (Duan et al., 2016; Wang et al., 2016; Mishra et al., 2017). Note that both DPT-greedy and PreDeToR are different than algorithmic distillation (Laskin et al., 2022; Lu et al., 2023) as they do not distill an existing RL algorithm. moreover, in contrast to DPT-greedy which is trained to predict the optimal action, the PreDeToR is trained to predict the reward for each of the actions. This enables the PreDeToR (similar to DPT-greedy) to show to potentially emergent online and offline strategies at test time that automatically align with the task structure, resembling posterior sampling.
As we discussed in the Section˜1, in decision-making, RL, and imitation learning the transformer models are trained using autoregressive action prediction (Yang et al., 2023). Similar methods have also been used in Large language models (Vaswani et al., 2017; Roberts et al., 2019). One of the more notable examples is the Decision Transformers (abbreviated as DT) which utilizes a transformer to autoregressively model sequences of actions from offline experience data, conditioned on the achieved return (Chen et al., 2021; Janner et al., 2021). This approach has also been shown to be effective for multi-task settings (Lee et al., 2022), and multi-task imitation learning with transformers (Reed et al., 2022; Brohan et al., 2022; Shafiullah et al., 2022). However, the DT methods are not known to improve upon their in-context data, which is the main thrust of this paper (Brandfonbrener et al., 2022; Yang et al., 2022b).
@@ A.2.2 Details of Baselines
(1) Thomp: This baseline is the stochastic $A$-action bandit Thompson Sampling algorithm from Thompson (1933); Agrawal & Goyal (2012); Russo et al. (2018); Zhu & Tan (2020). We briefly describe the algorithm below: At every round $t$ and each action $a$, Thomp samples $\gamma_{m,t}(a)\sim\mathcal{N}(\widehat{\mu}_{m,t-1}(a),\sigma^{2}/N_{m,t-1}(a))$, where $N_{m,t-1}(a)$ is the number of times the action $a$ has been selected till t1t-1, and $\widehat{\mu}_{m,t-1}(a)=\frac{\sum_{s=1}^{t-1}\widehat{r}_{m,s}\mathbf{1}(I_{s}=a)}{N_{m,t-1}(a)}$ is the empirical mean. Then the action selected at round $t$ is $I_{t}=\mathop{\rm arg\,max}_{a}\gamma_{m,t}(a)$. Observe that Thomp is not a deterministic algorithm like UCB (Auer et al., 2002). So we choose Thomp as the weak demonstrator $\pi^{w}$ because it is more exploratory than UCB and also chooses the optimal action, $a_{m,*}$, a sufficiently large number of times. Thomp is a weak demonstrator as it does not have access to the feature set $\mathcal{X}$ for any task $m$.
@@ A.3 Empirical Study: Comparison against K-armed bandits and DPT
Baselines: In the K-armed bandits We implement the same baselines discussed in Section˜4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LinUCB. In the linear and non-linear setting, we compare against DPT instead of DPT-greedy.
Settings: In the K-armed bandit setting we consider $d=6$, and the arms as canonical vectors $\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{6}$. For each task $m$, we choose the hidden parameter $\bm{\theta}_{m,*}$ similar to the linear bandit setting discussed in Section˜5. Note that this results in a K-armed bandit setting. For the linear and non-linear setting comparison, we use the same setting as Section˜5, and 4.
@@
Experimental Result: We observe these outcomes in Figure˜7. In Figure˜7(a) the demonstrator $\pi^{w}$ is the Thomp algorithm. Note that there is no structure across arms now, and sampling one arm gives no information about other arms in a task. We observe that PreDeToR-$\tau$ performs similarly to the demonstrator Thomp, and also shows that incorporating exploration is a sound technique. Also, AD performs similarly to the demonstrator Thomp. Both DPT-greedy and PreDeToR fail to learn the latent structure across the tasks and therefore do not learn any exploration strategy.
In Figure˜7(b) we show the linear bandit setting discussed in Section˜A.2. We observe that PreDeToR (-$\tau$) matches the performance of DPT, and LinUCB. Note that DPT has access to the optimal action per task, and LinUCB is the optimal oracle algorithm that leverages the structure information.
In Figure˜7(c) we show the non-linear bandit setting discussed in Section˜A.2. Again we observe that PreDeToR (-$\tau$) matches the performance of DPT and has lower cumulative regret than AD and LinUCB which fails to perform well in this non-linear setting due to its algorithmic design.
@@ A.4 Empirical Study: Bilinear Bandits
Baselines: We again implement the same baselines discussed in Section˜4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, and Thomp. Note that we do not implement the LinUCB and MLinGreedy for the bilinear bandit setting. However, we now implement the LowOFUL (Jun et al., 2019) which is optimal in the bilinear bandit setting.
@@
[…]
@@
(a) Rank 11 $\mathbf{\Theta}_{m,*}$
(b) Rank 22 $\mathbf{\Theta}_{m,*}$
@@
Experimental Result: We observe these outcomes in Figure˜8. In Figure˜8(a) we experiment with rank $1$ hidden parameter $\mathbf{\Theta}_{m,*}$ and set horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=30$, and $d=5$. In Figure˜8(b) we experiment with rank $2$ hidden parameter $\mathbf{\Theta}_{m,*}$ and set horizon $n=20$, $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR has lower cumulative regret than DPT-greedy, AD and Thomp. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Note that for this small horizon setting the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. In contrast PreDeToR learns the correlation of rewards across tasks and can perform well. Observe from Figure˜8(a), and 8(b) that PreDeToR has lower regret than Thomp and matches LowOFUL. Also, in this low-data regime it is not enough for LowOFUL to learn the underlying $\mathbf{\Theta}_{m,*}$ with high precision. Hence, PreDeToR also has slightly lower regret than LowOFUL. Note that the main objective of AD is to match the performance of its demonstrator. Most importantly it shows that PreDeToR can exploit the underlying latent structure and reward correlation better than DPT-greedy, and AD.
@@ A.5 Empirical Study: Latent Bandits
Baselines: We again implement the same baselines discussed in Section˜4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LowOFUL. However, we now implement a special LowOFUL (stated in Section˜A.4) which has knowledge of the shared latent parameters $\mathbf{U}$, and $\mathbf{V}$. We call this the LowOFUL (oracle) algorithm. Therefore LowOFUL (oracle) has knowledge of the problem parameters in the latent bandit setting and hence the name. Again note that we do not implement the LinUCB and MLinGreedy for the latent bandit setting.
@@
(a) Rank 11 $\mathbf{Z}_{m,*}$
(b) Rank 22 $\mathbf{Z}_{m,*}$
(c) Rank 33 $\mathbf{Z}_{m,*}$
@@
Experimental Result: We observe these outcomes in Figure˜9. In Figure˜9(a) we experiment with rank $1$ hidden parameter $\bm{\theta}_{m,*}\bm{\theta}_{m,*}^{\top}$ and latent parameters $\mathbf{u}\mathbf{v}^{\top}$ shared across the tasks and set horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=30$, and $d=5$. In Figure˜9(b) we experiment with rank $2$ hidden parameter $\mathbf{\Theta}_{m,*}$, and latent parameters $\mathbf{U}\mathbf{V}^{\top}$ and set horizon $n=20$, $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. In Figure˜9(c) we experiment with rank $3$ hidden parameter $\mathbf{\Theta}_{m,*}$, and latent parameters $\mathbf{U}\mathbf{V}^{\top}$ and set horizon $n=20$, $M_{\text{pre}}=300000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD and Thomp. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Note that for this small horizon setting the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. In contrast PreDeToR (-$\tau$) learns the correlation of rewards across tasks and is able to perform well. Observe from Figure˜9(a), 9(b), and 9(c) that PreDeToR has lower regret than Thomp and has regret closer to LowOFUL (oracle)which has access to the problem-dependent parameters. Hence. LowOFUL (oracle) outperforms PreDeToR (-$\tau$) in this setting. This shows that PreDeToR is able to exploit the underlying latent structure and reward correlation better than DPT-greedy, and AD.
@@ A.6 Connection between PreDeToR and Linear Multivariate Gaussian Model
$\mathbf{D}_{A}(i,i)$ $=\begin{cases}\frac{\sigma^{2}}{N_{n}(a)},\text{ if }N_{n}(a)>0\\ =0,\text{ if }N_{n}(a)=0\end{cases}\vskip-10.00002pt$ (6)
where the reward noise being $\sigma^{2}$ sub-Gaussian is known. Finally define the estimated reward covariance matrix $\mathbf{S}_{A}\in\mathbb{R}^{A\times A}$ as $\mathbf{S}_{A}(a,a^{\prime})=\widehat{\mu}_{n}(a)\widehat{\mu}_{n}(a^{\prime})$, where $\widehat{\mu}_{n}(a)$ is the empirical mean of action $a$ estimated from the pretraining data. This matrix captures the reward correlation between the pairs of actions $a,a^{\prime}\in[A]$. Then the posterior average mean estimator $\widehat{\mu}\in\mathbb{R}^{A}$ over all tasks is given by the following lemma. The proof is given in Section˜B.1.
@@
We hypothesize that the PreDeToR is learning the reward correlation covariance matrix from the training data $\mathcal{H}_{\text{train}}$ and acting greedily on it. To test this hypothesis, we consider the greedy BayesPred algorithm that first estimates $\mathbf{S}_{A}$ from the pretraining data. It then uses the LMMSE estimator in ˜1 to calculate the posterior mean vector $\widehat{\mu}$, and then selects $I_{t}=\mathop{\rm arg\,max}_{a}\widehat{\mu}(a)$ at each round $t$. Note that BayesPred is a greedy algorithm that always selects the most rewarding action (exploitation) without any exploration of sub-optimal actions. Also the BayesPred is an LMMSE estimator that leverages the linear reward structure and estimates the reward covariance matrix, and therefore can be interpreted as a lower bound to the regret of PreDeToR. The hypothesis that BayesPred is a lower bound to PreDeToR is supported by Figure˜10. In Figure˜10 the reward covariance matrix for BayesPred is estimated from the $\mathcal{H}_{\text{train}}$ by first running the Thomp ($\pi^{w}$). Observe that the BayesPred has a lower cumulative regret than PreDeToR and almost matches the regret of PreDeToR towards the end of the horizon. Also note that LinUCB has lower cumulative regret towards the end of horizon as it leverages the linear structure and the feature of the actions in selecting the next action.